OpenStudy (anonymous):

How to find min and max of x|2x + 5|

OpenStudy (anonymous):

can't imagine there is a max right?

OpenStudy (anonymous):

The function isn't bounded above or below.

OpenStudy (anonymous):

likewise for min. plug in values that go to minus infinity, thing goes there as well

OpenStudy (anonymous):

IN my caclulator it says there is

OpenStudy (anonymous):

what alchemista said (somewhat more elegantly)

OpenStudy (anonymous):

Perhaps you are confusing local minimums with the function overall.

OpenStudy (anonymous):

then i guess there must be. ask the calculator what it is

OpenStudy (anonymous):

I know....For future reference though, do I use the x value or y value or both if I do get a min and max?

OpenStudy (anonymous):

My apologies, I just came from Algebra ll and I always do hands on, calculator is a bit obtuse to me

OpenStudy (anonymous):

Pre Calc man oh man

OpenStudy (anonymous):

For the local minimum, you should be finding the roots of the quadratic and then between the roots will lie the minimum in this case.

OpenStudy (anonymous):

ah i see

OpenStudy (anonymous):

also

OpenStudy (anonymous):

how to do you find if it is bounded above, bounded below, bonded on its domain

OpenStudy (anonymous):

example y= 32

OpenStudy (anonymous):

The constant function \(f(x) = 32\)?

OpenStudy (anonymous):

y = 32 is a horizontal line. or rather the graph is. only range values isi {32}

OpenStudy (anonymous):

but how can you tell if it is bounded, etc, or not?

OpenStudy (anonymous):

In that case its just bounded above and below by 32

OpenStudy (anonymous):

Its simple for all \(x\) we have \(32 \leq f(x) \leq 32\)

OpenStudy (anonymous):

how can you tell of this one y= 2-x^2

OpenStudy (anonymous):

what things do i look for to find if it is bounded or not

OpenStudy (anonymous):

Well there is a maximum for that function, when x is 0

OpenStudy (anonymous):

The maximum is 2, and the function diverges to negative infinity to the left and to the right.

OpenStudy (anonymous):

So its bounded above by 2, but it is not bounded below.

OpenStudy (anonymous):

ah i see. how are you doing this without technical help?

OpenStudy (anonymous):

alright. i just got a huge impediment; y = 2^x

OpenStudy (anonymous):

Lets talk about the previous examples first. In any case without calculus it is difficult to prove any of these things rigorously. However, lets take the previous example, which is fairly simple.

OpenStudy (anonymous):

\(x^2 \geq 0\) for a all \(x \in \mathbb{R}\) and 0 when \(x=0\). So we have \(f(x) = 2 - x^2\). It is clear that the function will be highest when \(x^2\) is lowest which is at 0.

OpenStudy (anonymous):

When x is 0, the functions value is 2.

OpenStudy (anonymous):

So the function is bounded above by 2. However, as x gets very large to the right or to the left, the function gets very becomes a very large negative number. As you approach infinity in either direction the function diverges to negative infinity.

OpenStudy (anonymous):

seemingly irrelevant question, if infinity goes right,up or down, is it still positive?

OpenStudy (anonymous):

so would 2^x be bounded to its domain?

OpenStudy (anonymous):

What do you mean bounded to its domain?

OpenStudy (anonymous):

As for your previous question, a function is said to diverge to positive infinity, if the value becomes a larger and larger \(\textbf{positive}\) number and diverges to negative infinity if the value becomes and larger and larger \(\textbf{negative}\) number.

OpenStudy (anonymous):

sorry to deviate but my graph starts near the x-axis starting from the left curving up to infinity. fory y=2^x

OpenStudy (anonymous):

Well first of all, clearly \(2^x\) does not have a maximum. But it also does not have a minimum. Why do you think that is?

OpenStudy (anonymous):

It is certainly bounded below by zero. In fact zero is also the greatest lower bound. However, it does not acquire the value 0.

OpenStudy (anonymous):

it doenst have a place on the y axis? such as 2 - 3x or maybe i need to be rectified

OpenStudy (anonymous):

Did they talk about asymptotes in your class yet?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So it has a horizontal asymptote at 0

OpenStudy (anonymous):

Thanks you.